Solve the following exponential equation.
2.7^x=(0.6)^-x

Answer by Alan3354(69443)

2.7^x=(0.6)^-x
2.7^x=(3/5)^-x
2.7^x=(5/3)^x
x*log(2.7) = x*log(5/3)
x*(log(2.7) - log(5/3)) = 0
-------
x = 0